OpenStudy (anonymous):

If f(x)=3x-2 and g(x)=6-4x find f(x) + g(x) A 7+4x B4-x C-x-4 D-4-7x

OpenStudy (australopithecus):

just add the two functions

OpenStudy (australopithecus):

if I say y = a + b and x = m + c then, y + x = (m + c) + (a + b)

OpenStudy (anonymous):